OpenStudy (m):

M(T) = 88 st.dev.(T) = 8 R = -(5/8)T + 135 find st.dev.(R)

OpenStudy (anonymous):

Var(R)=(25/64)*Var(T)=25/64*64=25 st.dev(R)=sqrt(Var(R))=5